# Conceptualization of this study, Methodology, Writing

### v1

@@ 4 Experiments and Analysis
Table 4: Comparison with state-of-the-art models Across Datasets (RARP, Cholec80, JIGSAW)
Dataset Model Acc Edit F1@10 F1@25 F1@50 RARP MS-TCN Li et al. (2020) 69.0 - 70.7 ActionCLIP Psychogyios et al. (2023) 80.4 80.6 MA-TCN Van Amsterdam et al. (2022) 83.4 81.6 81.7 ASRF Ishikawa et al. (2021) $87.5\pm 3.7$ $83.6\pm 1.9$ $80.1\pm 3.2$ $70.8\pm 6.2$ $59.6\pm 3.2$ NETE Yi et al. (2022) $87.8\pm 3.6$ $83.9\pm 1.5$ $78.7\pm 3.6$ $68.5\pm 3.2$ $54.6\pm 2.4$ Asformer Yi et al. (2021) $\boldsymbol{88.8\pm 5.4}$ $\boldsymbol{84.1\pm 3.1}$ $\boldsymbol{84.3\pm 5.3}$ $68.1\pm 3.4$ $59.4\pm 3.1$ MSBATN(Ours) $88.7\pm 3.4$ $84.3\pm 2.7$ $84.0\pm 3.7$ $\boldsymbol{72.4\pm 2.6}$ $\boldsymbol{63.8\pm 1.5}$ Cholec80 ResNet Yi et al. (2022) 78.3 57.8 52.2 PhaseLSTM Neil et al. (2016) 80.7 61.2 62.4 EndoNet Twinanda et al. (2016) 81.9 TeCNO Czempiel et al. (2020) 88.6 ASRF Ishikawa et al. (2021) $82.5\pm 4.5$ $78.5\pm 3.5$ $84.3\pm 3.4$ $70.6\pm 5.7$ $60.1\pm 3.4$ Asformer Yi et al. (2021) $82.8\pm 4.7$ $74.5\pm 6.7$ $84.1\pm 2.4$ $70.8\pm 6.2$ $57.1\pm 3.2$ NETE Yi et al. (2022) $\boldsymbol{92.8\pm 5.0}$ $\boldsymbol{78.7\pm 9.4}$ $82.4\pm 3.4$ $70.1\pm 3.2$ $53.1\pm 4.8$ MSBATN(Ours) $89.4\pm 2.1$ $77.5\pm 8.0$ $\boldsymbol{84.4\pm 4.4}$ $\boldsymbol{73.1\pm 5.2}$ $\boldsymbol{60.2\pm 8.4}$ JIGSAW K-TCN Van Amsterdam et al. (2022) 83.8 86.3 90.4 MA-TCN Van Amsterdam et al. (2022) 86.8 90.1 93.6 ASRF Ishikawa et al. (2021) $87.8\pm 4.5$ $89.0\pm 2.5$ $80.4\pm 6.5$ $70.1\pm 3.6$ $69.8\pm 2.5$ NETE Yi et al. (2022) $90.8\pm 2.1$ $86.5\pm 3.5$ $74.4\pm 7.8$ $65.1\pm 3.7$ $58.6\pm 2.9$ Asformer Yi et al. (2021) $\boldsymbol{92.8\pm 3.3}$ $\boldsymbol{89.8\pm 6.7}$ $83.5\pm 4.2$ $70.1\pm 3.5$ $65.4\pm 1.8$ MSBATN(Ours) $86.1\pm 6.3$ $89.5\pm 6.5$ $\boldsymbol{84.4\pm 5.5}$ $\boldsymbol{76.1\pm 3.1}$ $\boldsymbol{70.7\pm 2.8}$
@@
We extracted every frame (256x 256x3) from the video sequences and used the Dinov2-Reg feature extractor to transform these frames into a 2048-dimensional feature space. These features serve as inputs to our encoder-decoder architecture. The encoder ingests inputs of shape (T, 2048) and processes them through a dilated acausal TCN, effectively reducing the dimensionality to (T,256) while capturing temporal dependencies over varying scales. These embeddings are further refined through ten layers of multi-head attention mechanisms (with 8 heads), incorporating advanced techniques such as sliding window attention for local dependencies and hierarchical attention for multi-scale temporal modelling. This combination enables the encoder to learn both fine-grained and long-range temporal patterns. The encoder produces frame-level predictions for action classes and boundary classes. To refine these initial predictions, three sequential decoders Zhang et al. (2023); Yi et al. (2021); Ishikawa et al. (2021); Yi et al. (2022) are employed. Each decoder shares a similar structure to the encoder but focuses on progressively refining predictions with higher accuracy and temporal smoothness. For optimization, we utilized the Adam optimizer along with the softmax activation function to ensure stable training. In contrast to some existing approaches Yi et al. (2021), positional encoding was omitted, as the hierarchical and sliding attention mechanisms sufficiently capture temporal ordering and dependencies. Our boundary-aware loss function is designed to enhance the model’s segmentation and transition detection capabilities. It combines a weighted Dice loss, which penalizes the model based on Intersection over Union (IoU), with a Gaussian Cosine Similarity loss that emphasizes smooth transitions by modelling centre-to-boundary similarity. This composite loss function ensures robust learning of temporal dynamics and accurate identification of action boundaries. We split the dataset into 80-20 ratios for training testing for Jigsaw and Cholec80. We used dropout and early stopping to handle over-fitting and set the max epoch to 120. We have considered three different datasets such as SAR-RARP50Psychogyios et al. (2023), JIGSAWSGao et al. (2014), Cholec80 Twinanda et al. (2016), and for more details, please refer to supplementary material.
4.1 Evaluation metrics
$\text{Recall}=\frac{\text{TP}_{IoU>T}}{\text{TP}_{IoU>T}+\text{FN}_{IoU>T}}$
Compute Segmental F1 Score:
[…]
The Edit Score assesses the sequence of predicted action segments by comparing it to the ground truth sequence. It is calculated based on the Levenshtein distance, which quantifies the minimum number of edit operations (insertions, deletions, substitutions) required to transform the predicted sequence into the ground truth sequence. A higher Edit Score indicates a closer match between the predicted and actual sequences, reflecting the model’s proficiency in capturing the correct order and transitions of actions. This metric is less sensitive to minor temporal shifts but penalizes over-segmentation and out-of-order predictions. Calculate the Levenshtein distance $D$ between the predicted and ground truth sequences. The Edit Score is then computed as Edit $\text{Edit Score}=1-\frac{D}{\max(L_{p},L_{g})}$ where $L_{p}$ and $L_{g}$ are the lengths of the predicted and ground truth sequences, respectively. The Segmental F1 Score is sensitive to over-segmentation and under-segmentation errors, penalizing models that divide actions into too many segments or fewer segments. In contrast, the Edit Score penalizes incorrect action order and over-segmentation but is more tolerant of slight temporal misalignments. The Segmental F1 Score evaluates the spatial and temporal precision of action boundaries, whereas the Edit Score assesses the structural and sequential accuracy of the predicted action sequence.
4.2 Results Analysis
Our Multi-staged Baundary aware network has three main differences form Asformer, First we employed acausal convolution layer in the initial layers of encoders and decoders. Secondly, we used hierarchical sparse multi-head window attention. Finally, we calculated centre sampled Boundary Aware Loss leverages the strengths of both cosine similarity and Gaussian functions, making it effective for segmentation tasks. We conducted the following experiments to demonstrate the effects of Causal and Acausal convolution along with hierarchical dual or single sliding window attentions.
4.2.1 Effect of Convolution and sliding window attention
We assessed the effectiveness of dilated causal and acausal convolutions, along with various attention settings during training. As illustrated in Table 3, the causal settings include: - Causal Self-Attention (CSAtt) - Causal Hierarchical Attention (CHAtt) - Causal Sliding Fixed Window with No Overlap (CSFWNO) - Causal Sliding Single Window with Overlap (CSSWO) - Causal Dual Sliding Window with Overlap (CDWSO) - Causal Hierarchical Fixed Window with Overlap (CHFWO) - Causal Hierarchical Dual Sliding Window with Overlap (CHDSWO)
In addition, we incorporated acausal convolutions, which include: - Acausal Self-Attention (ACSAtt) - Acausal Hierarchical Attention (ACHAtt) - Acausal Sliding Fixed Window with No Overlap (ACSFWNO) - Acausal Sliding Single Window with Overlap (ACSSWO) - Acausal Dual Sliding Window with Overlap (ACDWSO) - Acausal Hierarchical Fixed Window with Overlap (ACHFWO) - Acausal Hierarchical Dual Sliding Window with Overlap (ACHDSWO) Through ongoing evaluation of these settings, we concluded that for longer sequences, Dual Sliding Overlapping Window Attention outperformed Single Sliding Overlapping Window. Conversely, for shorter action sequences in the Jigsaw dataset, the Single Sliding Window demonstrated superior performance.
Table 3 highlights the effectiveness of proposed hierarchical sliding window attention with causal and acausal convolution. Settings with acausal convolution performed better than causal convolution as causal convolution restricts information flow to the present and past, ensuring that no future data influences the current output. In contrast, acausal convolution permits the output to depend on past, present, and future input values, which is useful during training. We also observe that self-attention and fixed hierarchical attention performed poorly. Vanilla self-attention computes attention scores between all token pairs, leading to a quadratic complexity. Such complexity becomes prohibitive for long sequences, limiting scalability. Hierarchical attention processes input sequences at multiple hierarchical levels with predetermined structures, which may not adapt well to varying local contexts within the data. Hierarchical Sliding Window Sparse Attention focuses on local neighbourhoods dynamically, capturing essential local patterns and structures. Stacking layers of such windowed attention results in a large receptive field, where top layers have access to all input locations and can build representations that incorporate information across the entire input. Single-Sliding Window Attention performs well on the Jigsaw dataset used in surgical training, where samples are shorter and backgrounds are simpler than in real-world datasets like SAR-RARP and Cholec80. In contrast, Dual-Sliding Window Attention performs better in SAR-RARP and Cholec80, which feature complex and dynamic actions. Single-Sliding Window Attention captures short-range dependencies effectively but may struggle with long-range connections, leading to a fragmented understanding in more complex videos. Meanwhile, Dual-Sliding Window Attention is better at capturing both short- and long-range dependencies, making it suitable for handling diverse action durations and transitions.
@@ 4.2.2 Effect of Start-End Boundary Aware Loss Function
Traditional loss functions, such as Binary Cross-Entropy (BCE), predominantly focus on frame-wise classification, which can result in either over-segmentation or under-segmentation. This approach has inherent limitations, particularly when addressing complex or ambiguous boundaries. Additionally, binary boundary classification may lead to instability during training due to substantial variations in loss values both within and between classes. The results of the proposed boundary-aware loss function show that our proposed approach effectively handled these challenges by capitalizing on the distribution of similarity scores between frames generated by decoders. Table 5 demonstrates the effectiveness of our method, which incorporates actual convolution and sliding window attention settings, delivering improved results across the five metrics outlined in the previous section compared to the binary-boundary-based approach.
@@ 5 Conclusion and Feature Work
We found that selecting the right attention module can be difficult, and neighbouring frames often lead to ambiguous results, causing over- and under-segmentation in frame-wise segmentation. Segment-wise refinement is more effective in improving understanding of surgical activities. Additionally, binary label boundaries are often inadequate because the start, centre, and end of actions vary by individual expertise, leading to poor performance. Our proposed framework addresses these issues with hierarchical sliding window attention for global-local context and includes a novel Boundary Segmentation Loss to better recognize central and boundary points. Our boundary-aware segmentation approach outperforms traditional binary detection methods. Experimental results show that our model matches or exceeds the performance of current state-of-the-art methods across various metrics, ensuring robustness in complex surgical environments.

### v2

@@ 4.2 Results
4.3 Ablation studies
We conducted an ablation study to demonstrate the effectiveness of the components in our architecture. Our configuration, which includes Acausal Convolution with Hierarchical Dual Sliding with Overlap, as shown in Table 3, outperforms other settings in long action segmentation. Additionally, we evaluated the computational complexity, as detailed in Table 5. Furthermore, we compared the effectiveness of the loss functions, as presented in Table 4.
4.3.1 Effect of Convolution and sliding window attention and overlapping
We assessed the effectiveness of convolutions, along with various attention settings during training. As illustrated in Table 3, the causal settings include: - Causal Self-Attention (CSAtt) - Causal Hierarchical Attention (CHAtt) - Causal Fixed Sliding Window with No Overlap (CFSWNO) - Causal Single Sliding Window with Overlap (CSSWO) - Causal Dual Sliding Window with Overlap (CDWSO) - Causal Hierarchical Fixed Window with Overlap (CHFWO) - Causal Hierarchical Dual Sliding Window with Overlap (CHDSWO) In addition, we incorporated acausal convolutions, which include: - Acausal Self-Attention (ACSAtt) - Acausal Hierarchical Attention (ACHAtt) - Acausal Fixed Sliding Window with No Overlap (ACFSWNO) - Acausal Single Sliding Window with Overlap (ACSSWO) - Acausal Dual Sliding Window with Overlap (ACDWSO) - Acausal Hierarchical Fixed Window with Overlap (ACHFWO) - Acausal Hierarchical Dual Sliding Window with Overlap (ACHDSWO)
Table 3 highlights the effectiveness of proposed hierarchical sliding window attention with causal and acausal convolution. Settings with acausal convolution performed better than causal convolution, as causal convolution restricts information flow to the present and past, ensuring that no future data influences the current output. In contrast, acausal convolution permits the output to depend on past, present, and future input values, which is useful during training. We also observe that self-attention and fixed hierarchical attention performed poorly. Vanilla self-attention computes attention scores between all token pairs, leading to a quadratic complexity. Such complexity becomes prohibitive for long sequences, limiting scalability. Hierarchical attention processes input sequences at multiple hierarchical levels with predetermined structures, which may not adapt well to varying local contexts within the data. Hierarchical Sliding Window Sparse Attention focuses on local neighbourhoods dynamically, capturing essential local patterns and structures. Stacking layers of such windowed attention results in a large receptive field, where the top layers have access to all input locations and can build representations that incorporate information across the entire input. Single-Sliding Window Attention performs well on the Jigsaw dataset used in surgical training, where samples are shorter and backgrounds are simpler than in real-world datasets like SAR-RARP and Cholec80. In contrast, Dual-Sliding Window Attention performs better in SAR-RARP and Cholec80, which feature complex and dynamic actions. Single-Sliding Window Attention captures short-range dependencies effectively but may struggle with long-range connections, leading to a fragmented understanding in more complex videos. Meanwhile, Dual-Sliding Window Attention is better at capturing both short- and long-range dependencies, making it suitable for handling diverse action durations and transitions.
@@ 4.3.1 Effect of Convolution and sliding window attention and overlapping
Table 5: Impact of HSWA Components (DSWA and HTA) and GFLOPs and parameters.
Model Acc F1@10 GFLOPs Parameters
Transformer Vaswani et al. (2017) 65.3 57.8 204.47 20.5181 M
ASRF Ishikawa et al. (2021) 87.6 80.3 136.42 13.6686 M
NETE Yi et al. (2022) 87.9 79.4 131.10 13.1454 M
Asformer Yi et al. (2021) 88.5 84.5 101.03 10.0027 M
DSWA 86.8 82.6 101.31 7.9750 M
HTA 84.4 76.2 76.78 7.8594 M
HSWA(DSWA+HTA) 88.7 84.5 104.53 11.9450 M
@@ 4.3.2 Effect of Start-End Boundary Aware Loss Function
Traditional loss functions, such as Binary Cross-Entropy (BCE), predominantly focus on frame-wise classification, which can result in either over-segmentation or under-segmentation. This approach has inherent limitations, particularly when addressing complex or ambiguous boundaries. Additionally, binary boundary classification may lead to instability during training due to substantial variations in loss values both within and between classes. The results of the proposed boundary-aware loss function show that our proposed approach effectively handled these challenges by capitalising on the distribution of similarity scores between frames generated by decoders. Table 4 demonstrates the effectiveness of our method, which incorporates acausal convolution and sliding window attention settings, delivering improved results across the five metrics outlined in the previous section compared to the binary-boundary-based approach.
@@ 5 Conclusion and Feature Work
We found that selecting the right attention module can be difficult, and neighbouring frames often lead to ambiguous results, causing over- and under-segmentation in frame-wise segmentation. Segment-wise refinement is more effective in improving the understanding of surgical activities. Additionally, binary label boundaries are often inadequate because the start, centre, and end of actions vary by individual expertise, leading to poor performance. Our proposed framework addresses these issues with hierarchical sliding window attention for global-local context and includes a novel Boundary Aware Segmentation Loss to better recognise central and boundary points. Our boundary-aware segmentation approach outperforms traditional binary detection methods. Experimental results show that our model matches or exceeds the performance of current state-of-the-art methods across various metrics, ensuring robustness in complex surgical environments.
@@
[…]
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the author(s) used Grammarly and ChatGPT in order to refine the writing, check grammar, and improve overall readability. After using these tools, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.